Lana ordered concert tickets that cost $7.50 for children and $12.00 for adults. She ordered 8 more children's tickets than adults' tickets. Her total bill was $138. How many of each type of ticket did she order?
Answer by mananth(16949)

Adult tickets x numbers
Child tickets y numbers
y-x=8
-1 x + 1 y = 8 .............1
12 x + 7.5 y = 138 .............2
Eliminate y
multiply (1)by -7.5
Multiply (2) by 1
7.5 x -7.5 y = -60
12 x 7.5 y = 138
Add the two equations
19.5 x = 78
/ 19.5
x = 4
plug value of x in (1)
-1 x + 1 y = 8
-4 + 1 y = 8
1 y = 8 + 4
1 y = 12
y = 12
Adult tickets 4 numbers
Child tickets 12 numbers
m.ananth@hotmail.ca
If you want to solve with one variable
let adults tickets be x numbers
children 's tickets = x+8
12*x+7.5(x+8) = 138
solve for x